OpenStudy (anonymous):

Find the value of the expression 4a^4 − 2b^2 + 40 when a = 2 and b = 7. Can someone Help plss

OpenStudy (tanya123):

Given: \(\ \sf\color{red}{a \color{black}{=} 2} \) \(\ \sf\color{blue}{b \color{black}{=} 7}\) Equation: \(\ \sf 4\color{red}{a}^4 − 2\color{blue}{b}^2 + 40\) Substitute the values you're given into your equation. Continue to solve.

OpenStudy (anonymous):

i already tried to solve this problem many times & i keep getting the same answer that is incorrect.. -.- @tanya123

OpenStudy (tanya123):

\(\ 4(\sf\color{red}{2})^2 - 2(\color{blue}{7})^2 + 40 \implies 4(4) - 2(49) + 40 = 16 - 98 + 40 = -42 \)

OpenStudy (anonymous):

its not one of the answer choices..

OpenStudy (tanya123):

Oh flutter, made a typo. 4a^4 =/= 4a^2 You get the idea. Try it yourself.

OpenStudy (tanya123):

Overall answer is 6, if you must know.

OpenStudy (anonymous):

i kind of get it now.. but thank you :)

OpenStudy (tanya123):

You're welcome.